# LASSI: An LLM-based Automated Self-Correcting Pipeline for Translating Parallel Scientific Codes

### v1

@@ Abstract
This paper addresses the problem of providing a novel approach to sourcing significant training data for LLMs focused on science and engineering. In particular, a crucial challenge is sourcing parallel scientific codes in the ranges of millions to billions of codes. To tackle this problem, we propose an automated pipeline framework, called LASSI, designed to translate between parallel programming languages by bootstrapping existing closed- or open-source LLMs. LASSI incorporates autonomous enhancement through self-correcting loops where errors encountered during compilation and execution of generated code are fed back to the LLM through guided prompting for debugging and refactoring. We highlight the bi-directional translation of existing GPU benchmarks between OpenMP target offload and CUDA to validate LASSI.
The results of evaluating LASSI with different application codes across four LLMs demonstrate the effectiveness of LASSI for generating executable parallel codes, with 80% of OpenMP to CUDA translations and 85% of CUDA to OpenMP translations producing the expected output. We also observe approximately 78% of OpenMP to CUDA translations and 62% of CUDA to OpenMP translations execute within 10% of or at a faster runtime than the original benchmark code in the same language.
@@ I Introduction
The problem addressed in this paper is that of providing a novel approach to sourcing significant training data for Large Language Models (LLMs) focused on science and engineering, a key objective of the Trillion Parameter Consortium (TPC) [1]. TPC brings together international communities that encompass three areas: (1) those working to advance AI methods with a focus on LLMs, (2) those with existing or emerging exascale platforms necessary for training LLMs, and (3) those who will use the resulting LLMs to address problems in science and engineering. In particular, this paper is focused on the need to source parallel scientific codes.
To adequately source parallel scientific codes, it is important to have a framework that can easily generate millions to billions of codes in different programming languages widely used in the sciences, such as FORTRAN, C++, CUDA, HIP, OpenMP, Julia, and SYCL. Further, the generated code should be “good” in the sense that it should compile, execute, be performant, and correct. The framework should be automated to facilitate the large number of codes needed. To address this problem, we present LASSI, an LLM-based Automated Self-correcting pipeline for generating parallel ScIentific codes. Currently, LASSI features a pipeline that automatically generates code, test for compilation, and checks execution. Future work will consider the correctness and performance.
In this paper, LASSI is used to translate codes between OpenMP and CUDA executed on an NVIDIA A100 GPU. For the case of OpenMP code, we use the offload to GPU feature. We provide the results of using four LLMs with ten codes from the HeCBench suite [2], which serve as a basis for comparison of the execution time of the code generated by LASSI. We observe the importance of incorporating the capability within the pipeline to provide feedback from compilation and execution errors for self-correcting. Further, we observe that approximately 78% of the translations from OpenMP to CUDA and 62% of the translations from CUDA to OpenMP are within 10% of or faster than the original codes in HeCBench in the same programming language.
@@ II Related Work
Many existing commercial and open-source LLMs, such as GPT-4 [3], Codestral [4], StarCoder [5], and Code Llama [6], are trained on massive collections of shared code primarily developed in widely-used programming languages like Python and JavaScript. Specialized open-source models are being released frequently that are further tuned to general purpose coding tasks, e.g., Wizard Coder [7] and DeepSeek Coder v2 [8]. However, these state-of-the-art code-centric models still lack sufficient training data for parallel scientific codes, especially those typically utilized for HPC scientific applications. This gap may stem from the limited volume of scientific developers who actively share code in the HPC domain, compared to the broader industries of web and mobile app development, which typically do not require the use of HPC for code execution.
LLMs routinely demonstrate highly effective general capabilities when incorporating context, such as domain-specific knowledge [9] or even an entertaining personality type [10], which guides generated responses specific to the context. The learned representations of human language by an LLM are harnessed along with this context to reduce so-called hallucination. The addition of context involves enhancing the probabilities during inference towards responses that include the provided source content over other content learned in the weights of the model. The technique of retrieval augmented generation (RAG) [11] leverages this LLM behavior. We adopt a similar yet simplified approach, aiming to enhance an LLM’s performance in parallel scientific code translation by employing carefully crafted prompts imbued with contextual knowledge and tailored programming expectations.
@@ III LASSI: Automated Self-Correcting Pipeline
We propose LASSI, an automated pipeline framework designed to translate between parallel programming languages by bootstrapping existing closed- or open-source LLMs. LASSI incorporates domain knowledge as a core feature, offering the advantage of tailoring prompts that guide the LLM towards synthesizing desired programming languages and performance outcomes. This is particularly beneficial given that the model does not have a high-quality foundation training in parallel coding techniques. Furthermore, the impact prompting has on the quality of an LLM response is significant. The prompting strategies and techniques presented here suggest reasonable performance, which were developed through extensive trial-and-error.
With a pipeline that is LLM-agnostic, we acknowledge that continuous effort is required to optimize prompt content, especially as new LLMs are released in the future. An intriguing avenue towards this end is to leverage an LLM to help design its prompts [19], an approach we explored and incorporate into our solution to improve generated code results.
Figure 1 summarizes the LASSI architecture with the LLM at the core of all operations, taking input from an extensive prompting strategy with domain knowledge and feedback from compilation and execution errors that autonomously guide the generation of working code. In the following subsections, we provide a detailed description of LASSI, including the specific prompting utilized to guide an LLM to generate the reported results through automated self-correction iterations.
@@ III-A Source Code Preparation
The initial step of the LASSI framework establishes an experimental baseline by compiling and executing the original target language code. This step ensures the viability of our approach by providing a basis for comparison with the code generated by LASSI. Upon successful execution, the standard output of the executed code is captured for later comparison with the output of the generated code. This initial step also serves to verify that LASSI’s compilation command is appropriate for the local compute platform because the same command will be used for compiling the LASSI generated code. If an error occurs, LASSI halts and does not move forward with the translation until the code is corrected by the user.
LASSI also checks that the original source language code compiles and executes in the local environment before processing through the translation pipeline. If an error occurs, again LASSI halts and does not move forward with the translation until the code is corrected by the user.
@@ III-B Programming Language-Specific Context Preparation
A key feature of the LASSI prompting strategy is to incorporate specific knowledge tailored to the target programming language (CUDA or OpenMP). We integrated content sourced from the official CUDA manual for the translation in the OpenMP to CUDA pipeline, and content from OpenMP resources for the CUDA to OpenMP translations.
@@ III-D Self-Correcting Loops for Autonomous Improvement
LASSI incorporates a novel self-correction routine. The LLM attempts to rectify errors by re-prompting with the context of specific compile or execute errors. This unique integration provides some autonomous control within the code generation pipeline. In the following subsection, we detail how errors encountered during compilation and execution are returned to the LLM through guided prompting for debugging and refactoring the generated code.
@@ III-D1 Integrated Code Compilation with Self-Correction
New code is generated again by the LLM followed by another compilation attempt with a capture of any resulting error messages. This iteration continues until no error is returned when compiling the generated code.
@@ III-D2 Integrated Code Execution with Self-Correction
New code is generated once again by the LLM followed by a compilation attempt with a capture of resulting error messages. If a compile error occurs again, then the pipeline remains in the compilation self-correction loop. This iteration through the compiler and execution attempts continue until no error is returned from executing the generated code.
@@ IV Benchmark Codes and LLMs
The goal of LASSI is to translate existing science code from one parallel programming language to another. To accomplish this task we leverage the HeCBench repository [2] for our code base. HeCBench offers an extensive curation of open-source heterogeneous computing applications available in OpenMP, CUDA, HPI, and SYCL. For this study, we focus on bi-directional translation between GPU benchmarking codes written in OpenMP with target offload and CUDA. We selected a suite of HeCBench codes to use the application version in one language as the source for our pipeline to translate to another language and the corresponding version of the same application in the target language to compare with the LASSI generated code. Moreover, we ensured diversity in computational categories to demonstrate the robustness in the translation capabilities. Following HeCBench’s categorization, we selected ten applications across nine categories for our test cases, as listed in Table IV.
We compiled and executed each HeCBench test case written in either CUDA or OpenMP using the same compilers and flags. Further, identical input parameters were used with LASSI for execution on the same compute resources. The runtimes were measured for each benchmark code to compare the runtime for the LASSI generated code. These runtimes are also listed in Table IV and represent an average runtime of three executions on an NVIDIA A100 GPU. The average is used because the standard deviation is small due to the single-user access to this local server.
@@
Panel A: GPT-4 Large and Codestral 22B 8-bit LLMs GPT-4 Codestral Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr matrix-rotate 1.2039 1.0333 0.44 0.83 1 1.0398 1.1964 0.31 0.68 0 jacobi 0.6746 1.2809 0.63 0.52 0 0.3395 2.5452 0.54 0.47 0 layout 0.6983 0.5854 0.63 0.68 0 0.4045 1.0106 0.50 0.45 0 atomicCost 45.8775 0.5854 0.63 0.68 0 12.0574 3.6425 0.58 0.50 0 dense-embedding N/A N/A N/A N/A N/A 0.8823 0.9130 0.49 0.34 1 pathfinder 0.6306 0.8595 0.50 0.36 0 0.2677 2.0246 0.39 0.18 1 bsearch N/A N/A N/A N/A N/A 0.2878 1.1372 0.29 0.22 0 entropy 0.5885 4.0596 0.64 0.57 1 3.9575 0.6037 0.37 0.24 2 colorwheel 0.3271 0.9199 0.70 0.51 3 N/A N/A N/A N/A N/A randomAccess N/A N/A N/A N/A N/A 8.8905 0.5640 0.67 0.55 2
Panel B: Wizard Coder 33B 8-bit and DeepSeek Coder v2 16B F16 LLMs Wizard Coder DeepSeek Coder v2 Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr matrix-rotate 1.1404 1.0909 0.37 0.61 0 1.0808 1.1510 0.32 0.64 0 jacobi 0.2892 2.9879 0.31 0.28 0 0.8327 1.0377 0.44 0.21 1 layout 0.4055 1.0081 0.53 0.53 0 0.6433 0.6355 0.46 0.51 0 atomicCost 116.2879 0.3777 0.59 0.57 0 93.1467 0.4715 0.58 0.47 1 dense-embedding 0.8137 0.9899 0.64 0.54 0 N/A N/A N/A N/A N/A pathfinder 0.4804 1.1282 0.47 0.39 0 0.6821 0.7946 0.33 0.22 0 bsearch 0.2706 1.2095 0.35 0.32 1 0.2675 1.2236 0.42 0.41 0 entropy 2.3551 1.0144 0.50 0.42 0 2.4239 0.9856 0.58 0.54 0 colorwheel 0.2997 1.0040 0.64 0.41 2 N/A N/A N/A N/A N/A randomAccess N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
@@
Panel A: GPT-4 Large and Codestral 22B 8-bit LLMs GPT-4 Codestral Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr matrix-rotate 1.0857 1.0869 0.80 0.93 0 1.0398 1.1349 0.76 0.90 0 jacobi 42.8133 1.3392 0.45 0.43 0 N/A N/A N/A N/A N/A layout 0.2755 0.9339 0.60 0.67 0 0.4040 0.6369 0.43 0.51 1 atomicCost 219.5494 0.2055 0.84 0.80 0 72.0812 0.6260 0.77 0.66 0 dense-embedding N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A pathfinder 0.2416 3.0033 0.40 0.27 1 0.2659 2.7288 0.14 0.09 34 bsearch 0.0045 3.1111 0.41 0.37 0 0.2811 0.0498 0.47 0.57 0 entropy 1.4200 2.4392 0.65 0.46 1 3.9527 0.8763 0.71 0.70 0 colorwheel 0.0044 0.7273 0.87 0.74 0 0.0023 1.3913 0.79 0.81 0 randomAccess 7.9183 0.9997 0.85 0.83 0 8.8873 0.8907 0.65 0.75 0
Panel B: Wizard Coder 33B 8-bit and DeepSeek Coder v2 16B F16 LLMs Wizard Coder DeepSeek Coder v2 Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr matrix-rotate 0.7645 1.5435 0.44 0.51 2 11.0047 0.1072 0.58 0.80 0 jacobi 1.4433 39.7252 0.42 0.43 4 1.6659 34.4171 0.37 0.28 1 layout 0.1326 1.9404 0.19 0.54 0 0.1639 1.5699 0.19 0.47 2 atomicCost 35.8374 1.2591 0.37 0.23 1 0.6805 66.3104 0.54 0.46 1 dense-embedding 56.6443 1.0090 0.54 0.44 0 N/A N/A N/A N/A N/A pathfinder 0.3914 1.8539 0.26 0.15 0 N/A N/A N/A N/A N/A bsearch 0.0158 0.8861 0.37 0.41 1 0.0048 2.9167 0.38 0.42 2 entropy 3.9525 0.8763 0.70 0.60 0 7.8830 0.4394 0.63 0.48 1 colorwheel 0.0046 0.6957 0.67 0.44 1 0.0146 0.2192 0.73 0.63 2 randomAccess 8.8987 0.8896 0.59 0.49 1 N/A N/A N/A N/A N/A

### v2

@@ Abstract
This paper addresses the problem of providing a novel approach to sourcing significant training data for LLMs focused on science and engineering. In particular, a crucial challenge is sourcing parallel scientific codes in the ranges of millions to billions of codes. To tackle this problem, we propose an automated pipeline framework called LASSI, designed to translate between parallel programming languages by bootstrapping existing closed- or open-source LLMs. LASSI incorporates autonomous enhancement through self-correcting loops where errors encountered during the compilation and execution of generated code are fed back to the LLM through guided prompting for debugging and refactoring. We highlight the bi-directional translation of existing GPU benchmarks between OpenMP target offload and CUDA to validate LASSI. The results of evaluating LASSI with different application codes across four LLMs demonstrate the effectiveness of LASSI for generating executable parallel codes, with 80% of OpenMP to CUDA translations and 85% of CUDA to OpenMP translations producing the expected output. We also observe approximately 78% of OpenMP to CUDA translations and 62% of CUDA to OpenMP translations execute within 10% of or at a faster runtime than the original benchmark code in the same language.
@@ I Introduction
Existing large language models (LLMs) such as GPT-4 [1], Codestral [2], StarCoder [3], and Code Llama [6] trained on public datasets have demonstrated limitations in generating high-quality scientific code, underscoring the need for a novel approach to source adequate training data for a new LLM specializing in science code development. The problem addressed in this paper is that of providing a novel approach to sourcing significant training data for LLMs specialized on science and engineering, a key objective of the Trillion Parameter Consortium (TPC) [4]. TPC brings together international communities that encompass three areas: (1) those working to advance AI methods with a focus on LLMs, (2) those with existing or emerging exascale platforms necessary for training LLMs, and (3) those who will use the resulting LLMs to address problems in science and engineering. In particular, this paper focuses on the need to source parallel scientific codes.
To adequately source parallel scientific codes, it is important to have a framework that can easily generate millions to billions of codes in different programming languages widely used in the sciences, such as C++, C, FORTRAN, CUDA, HIP, OpenMP, Julia, and SYCL. Further, the generated code should be “good” in the sense that it should compile, execute, be performant, and correct. The framework should be automated to facilitate the large number of codes needed. To address this problem, we present LASSI, an LLM-based Automated Self-correcting pipeline for generating parallel ScIentific codes. Currently, LASSI features a pipeline that automatically generates code, tests for compilation, and checks execution. Future work will consider code performance and correctness.
In this paper, we demonstrate an implementation of LASSI to translate codes between OpenMP and CUDA executed on an NVIDIA A100 GPU. For the case of OpenMP code, we use the offload to GPU feature. We provide the results of using four LLMs with ten codes from the HeCBench suite [5], which serve as a basis for comparison of the execution time of the code generated by LASSI. We observe the importance of incorporating the capability within the pipeline to provide feedback from compilation and execution errors for self-correcting. Further, we observe that approximately 78% of the translations from OpenMP to CUDA and 62% of the translations from CUDA to OpenMP are within 10% of or faster than the original codes in HeCBench in the same programming language.
@@ II Related Work
Many existing commercial and open-source LLMs, such as GPT-4 [1], Codestral [2], StarCoder [3], and Code Llama [6], are trained on massive collections of shared code primarily developed in widely-used programming languages like Python and JavaScript. Specialized open-source models that are further tuned to general purpose coding tasks are being released frequently, e.g., Wizard Coder [7] and DeepSeek Coder v2 [8]. However, these state-of-the-art code-centric models still lack sufficient training data for parallel scientific codes, especially those typically utilized for HPC scientific applications. This gap may stem from the limited volume of scientific developers who actively share code in the HPC domain compared to the broader industries of web and mobile app development, which typically do not require HPC for code execution.
LLMs routinely demonstrate highly effective general capabilities when incorporating context, such as domain-specific knowledge [9, 10] or even an entertaining personality type [11], which guides generated responses specific to the context. The learned representations of human language by an LLM are harnessed along with this context to reduce so-called hallucination. The addition of context involves enhancing the probabilities during inference towards responses that include the provided source content over other content learned in the weights of the model. The technique of retrieval augmented generation (RAG) [12] leverages this LLM behavior. We adopt a similar yet simplified approach, aiming to enhance an LLM’s performance in parallel scientific code translation by employing carefully crafted prompts imbued with contextual knowledge and tailored programming expectations.
@@ III LASSI: Automated Self-Correcting Pipeline
We propose LASSI, an automated pipeline framework designed to translate between parallel programming languages by bootstrapping existing closed- or open-source LLMs. LASSI incorporates domain knowledge as a core feature, offering the advantage of tailoring prompts that guide the LLM toward synthesizing desired programming languages and performance outcomes. This is particularly beneficial given that the model does not have high-quality foundation training in parallel coding techniques. Furthermore, the impact prompting has on the quality of an LLM response is significant. The prompting strategies and techniques presented here suggest reasonable performance, which were developed through extensive trial and error.
With an LLM-agnostic pipeline, we acknowledge that continuous effort is required to optimize prompt content, especially as new LLMs are released in the future. An intriguing avenue towards this end is to leverage an LLM to help design its prompts [20], an approach we explored and incorporate into our solution to improve generated code results.
Figure 1 summarizes the LASSI architecture with the LLM at the core of all operations, taking input from an extensive prompting strategy with domain knowledge and feedback from compilation and execution errors that autonomously guide the generation of working code. In the following subsections, we provide a detailed description of LASSI, including the specific prompting utilized to guide an LLM in generating the reported results through automated self-correction iterations.
@@ III-A Source Code Preparation
The initial step of the LASSI framework establishes an experimental baseline by compiling and executing the original target language code. This step ensures the viability of our approach by providing a basis for comparison with the code generated by LASSI. Upon successful execution, the standard output of the executed code is captured for later comparison with the output of the generated code. This initial step also serves to verify that LASSI’s compilation command is appropriate for the local compute platform because the same command will be used to compile the LASSI-generated code. If an error occurs, LASSI halts and does not move forward with the translation until the user corrects the code.
LASSI also checks that the original source language code compiles and executes in the local environment before processing through the translation pipeline. If an error occurs, again, LASSI halts and does not move forward with the translation until the code is corrected by the user.
@@ III-B Programming Language-Specific Context Preparation
A key feature of the LASSI prompting strategy is the incorporation of specific knowledge tailored to the target programming language (CUDA or OpenMP). We integrated content sourced from the official CUDA manual for the translation in the OpenMP to CUDA pipeline and content from OpenMP resources for the CUDA to OpenMP translations.
@@ III-D Self-Correcting Loops for Autonomous Improvement
LASSI incorporates a novel self-correction routine. The LLM attempts to rectify errors by re-prompting along with the context of specific compile or execute errors. This unique integration provides some autonomous control within the code generation pipeline. In the following subsection, we detail how errors encountered during compilation and execution are returned to the LLM through guided prompting for debugging and refactoring the generated code.
@@ III-D1 Integrated Code Compilation with Self-Correction
New code is generated again by the LLM, followed by another compilation attempt with a capture of any resulting error messages. This iteration continues until no error is returned when compiling the generated code.
@@ III-D2 Integrated Code Execution with Self-Correction
New code is generated once again by the LLM, followed by a compilation attempt with a capture of resulting error messages. If a compile error occurs again, then the pipeline remains in the compilation self-correction loop. This iteration through the compiler and execution attempts continue until no error is returned from executing the generated code.
@@ IV Benchmark Codes and LLMs
The goal of LASSI is to translate existing science code from one parallel programming language to another. We leverage the HeCBench repository [5] for our code base to accomplish this task. HeCBench offers an extensive curation of open-source heterogeneous computing applications available in OpenMP, CUDA, HPI, and SYCL. Because HeCBench offers hundreds of parallel code examples refactored in multiple languages, this benchmark suite is practical for prototyping the LASSI pipeline by enabling direct comparisons between its generated code and the source code in the suite.
For this study, we focus on bi-directional translation between GPU benchmarking codes written in OpenMP with target offload and CUDA. We selected HeCBench codes to use the application version in one language as the source for our pipeline to translate to another language and the corresponding version of the same application in the target language to compare with the LASSI-generated code. Moreover, we ensured diversity in computational categories to demonstrate the robustness of the translation capabilities. Following HeCBench’s categorization, we selected ten applications across nine categories for our test cases, as listed in Table IV.
We compiled and executed each HeCBench test case written in either CUDA or OpenMP using the same compilers and flags. Further, identical input parameters were used with LASSI for execution on the same compute resources. The runtimes were measured for each benchmark code to compare the runtime for the LASSI-generated code. These runtimes are listed in Table IV and represent an average runtime of three executions on an NVIDIA A100 GPU. The average is used because the standard deviation is small due to the single-user access to this local server.
@@
(a) Panel A: GPT-4 Large and Codestral 22B 8-bit LLMs
GPT-4 Codestral
Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr
matrix-rotate 1.2039 1.0333 0.44 0.83 1 1.0398 1.1964 0.31 0.68 0
jacobi 0.6746 1.2809 0.63 0.52 0 0.3395 2.5452 0.54 0.47 0
layout 0.6983 0.5854 0.63 0.68 0 0.4045 1.0106 0.50 0.45 0
atomicCost 45.8775 0.5854 0.63 0.68 0 12.0574 3.6425 0.58 0.50 0
dense-embedding N/A N/A N/A N/A N/A 0.8823 0.9130 0.49 0.34 1
pathfinder 0.6306 0.8595 0.50 0.36 0 0.2677 2.0246 0.39 0.18 1
bsearch N/A N/A N/A N/A N/A 0.2878 1.1372 0.29 0.22 0
entropy 0.5885 4.0596 0.64 0.57 1 3.9575 0.6037 0.37 0.24 2
colorwheel 0.3271 0.9199 0.70 0.51 3 N/A N/A N/A N/A N/A
randomAccess N/A N/A N/A N/A N/A 8.8905 0.5640 0.67 0.55 2
(b) Panel B: Wizard Coder 33B 8-bit and DeepSeek Coder v2 16B F16 LLMs
Wizard Coder DeepSeek Coder v2
Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr
matrix-rotate 1.1404 1.0909 0.37 0.61 0 1.0808 1.1510 0.32 0.64 0
jacobi 0.2892 2.9879 0.31 0.28 0 0.8327 1.0377 0.44 0.21 1
layout 0.4055 1.0081 0.53 0.53 0 0.6433 0.6355 0.46 0.51 0
atomicCost 116.2879 0.3777 0.59 0.57 0 93.1467 0.4715 0.58 0.47 1
dense-embedding 0.8137 0.9899 0.64 0.54 0 N/A N/A N/A N/A N/A
pathfinder 0.4804 1.1282 0.47 0.39 0 0.6821 0.7946 0.33 0.22 0
bsearch 0.2706 1.2095 0.35 0.32 1 0.2675 1.2236 0.42 0.41 0
entropy 2.3551 1.0144 0.50 0.42 0 2.4239 0.9856 0.58 0.54 0
colorwheel 0.2997 1.0040 0.64 0.41 2 N/A N/A N/A N/A N/A
randomAccess N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
@@
(a) *Panel A: GPT-4 Large and Codestral 22B 8-bit LLMs
GPT-4 Codestral
Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr
matrix-rotate 1.0857 1.0869 0.80 0.93 0 1.0398 1.1349 0.76 0.90 0
jacobi 42.8133 1.3392 0.45 0.43 0 N/A N/A N/A N/A N/A
layout 0.2755 0.9339 0.60 0.67 0 0.4040 0.6369 0.43 0.51 1
atomicCost 219.5494 0.2055 0.84 0.80 0 72.0812 0.6260 0.77 0.66 0
dense-embedding N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
pathfinder 0.2416 3.0033 0.40 0.27 1 0.2659 2.7288 0.14 0.09 34
bsearch 0.0045 3.1111 0.41 0.37 0 0.2811 0.0498 0.47 0.57 0
entropy 1.4200 2.4392 0.65 0.46 1 3.9527 0.8763 0.71 0.70 0
colorwheel 0.0044 0.7273 0.87 0.74 0 0.0023 1.3913 0.79 0.81 0
randomAccess 7.9183 0.9997 0.85 0.83 0 8.8873 0.8907 0.65 0.75 0
(b) *Panel B: Wizard Coder 33B 8-bit and DeepSeek Coder v2 16B F16 LLMs
Wizard Coder DeepSeek Coder v2
Runtime (s) Ratio Sim-T Sim-L Self-corr Runtime (s) Ratio Sim-T Sim-L Self-corr
matrix-rotate 0.7645 1.5435 0.44 0.51 2 11.0047 0.1072 0.58 0.80 0
jacobi 1.4433 39.7252 0.42 0.43 4 1.6659 34.4171 0.37 0.28 1
layout 0.1326 1.9404 0.19 0.54 0 0.1639 1.5699 0.19 0.47 2
atomicCost 35.8374 1.2591 0.37 0.23 1 0.6805 66.3104 0.54 0.46 1
dense-embedding 56.6443 1.0090 0.54 0.44 0 N/A N/A N/A N/A N/A
pathfinder 0.3914 1.8539 0.26 0.15 0 N/A N/A N/A N/A N/A
bsearch 0.0158 0.8861 0.37 0.41 1 0.0048 2.9167 0.38 0.42 2
entropy 3.9525 0.8763 0.70 0.60 0 7.8830 0.4394 0.63 0.48 1
colorwheel 0.0046 0.6957 0.67 0.44 1 0.0146 0.2192 0.73 0.63 2
randomAccess 8.8987 0.8896 0.59 0.49 1 N/A N/A N/A N/A N/A